There were 38 tubes containing exactly 40 Smarties. The observed distribution of yellow beans in these 38 tubes is given in the following table.

Number of yellow beans 0 1 2 3 4 5 6 7 8 ≥9
Frequency 0 2 10 6 7 5 2 5 1 0

Calculate the proportions of tubes of 40 Smarties that contain 0, 1, 2, . . . , 8,≥ 9 yellow beans. Give the proportions rounded to three decimal places.
Answer #1

Given is the following frequency distribution of yellow beans in these 38 tubes ,

Number of yellow beans 0 1 2 3 4 5 6 7 8 ≥9
Frequency 0 2 10 6 7 5 2 5 1 0

To obtain the corresponding probability distribution of tubes of 40 Smarties that contain 0, 1, 2, . . . , 8,≥ 9 yellow beans,
and is obtained as follows,

Number of Yellow Beans Proportion of tubes of 40 Smarties
0 0
1 1/38 = 0.026
2 10/38 = 0.263
3 6/38 = 0.158
4 7/38 = 0.184
5 5/38 = 0.132
6 2/38 = 0.053
7 5/38 = 0.132
8 1/38 = 0.026
0

Answer : Above table is the required corresponding proportions of tube of 40 Smarties